# Another algorithm template

### v11

@@ 5 Elliptic PDEs with lognormal inputs
where $S_{\Lambda}u:=\sum_{{\boldsymbol{s}}\in\Lambda}u_{\boldsymbol{s}}$ for $\Lambda\subset{\mathbb{F}}$, $k_{n}:=\lfloor\log_{2}\xi_{n}\rfloor$ and for $k\in{\mathbb{N}}_{0}$ and $\xi>1$,
@@
where $k_{\xi}:=\big{\lfloor}\frac{1}{\alpha q_{1}}\log_{2}\xi\big{\rfloor}$ and $\Lambda_{k}(\xi)$ is defined as in (5.3), $\xi>1$. Moreover, $\Lambda_{k}(\xi)$ are downward closed sets, and consequently, the sequence $\big{\{}\Lambda_{k}(\xi)\big{\}}_{k=0}^{k_{\xi}}$ is nested in the inverse order, i.e., $\Lambda_{k^{\prime}}(\xi)\subset\Lambda_{k}(\xi)$ if $k^{\prime}>k$, and $\Lambda_{0}(\xi)$ is the largest and $\Lambda_{k_{\xi}}(\xi)$ is the smallest. Hence, the index set $G(\xi)$ can be constructed as in Algorithm 1.
@@ 28    $k\leftarrow k+1$
Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that for the interpolation operator ${\mathcal{I}}_{G(\xi_{n})}:{\mathcal{L}}_{2}^{\mathcal{E}}(W)\to{\mathcal{V}}(G(\xi_{n}))$, we have that $\dim{\mathcal{V}}(G(\xi_{n}))\leq n$ and
@@
The rate $\alpha$ corresponds to the spatial approximation of a single function in $W$ as given by (1.5). The rate $\beta$ is given by (3.21). The constant $C$ in (5.4) is independent of $u$ and $n$.
@@
Let $0<p\leq 2$. Under the hypothesis of Lemma 5.1 for the spaces $W^{1}=V$ with some $0<q_{1}=q<2$. For $\xi>1$, let $\Lambda(\xi)$ be the set defined by (3.39) for $\sigma_{\boldsymbol{s}}=\sigma_{1;{\boldsymbol{s}}}$ as in (5.1). Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda(\xi_{n}))|\leq n$ and
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
@@
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,{\rm d}\gamma({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}\langle\phi,v\rangle\right|\leq C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (5.7)
@@
$\beta:=\left(\frac{2}{q_{1}}-\frac{1}{2}\right)\frac{\alpha}{\alpha+\delta},\quad\delta:=\frac{2}{q_{1}}-\frac{2}{q_{2}}.$
@@
where $k_{n}:=\lfloor\log_{2}\xi_{n}\rfloor$ and for $k\in{\mathbb{N}}_{0}$ and $\xi>1$,
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
@@
    Let $\phi\in V^{\prime}$ a bounded linear functional on $V$. There exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}(\xi_{n})}\langle\phi,u\rangle\right|\ \leq\ Cn^{-(2/q-1/2)}.$ (5.9)
@@
Let Assumption I hold. Let the assumptions of Lemma 5.1 hold for the spaces $W^{1}=V$ and $W^{2}=W$ with some $0<q_{1}\leq q_{2}<\infty$ with $q_{1}<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. For $\xi>1$, let $G_{2}(\xi)$ be the set defined as in (4.9) for $\sigma_{r;{\boldsymbol{s}}}$ as in (5.1), $r=1,2$. Then we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G(\xi_{n}))\leq n$ and
@@
    Let $\phi\in V^{\prime}$ be a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G(\xi_{n}))\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{{\mathcal{Q}}}_{G_{2}(\xi_{n})}\langle\phi,v\rangle\right|\leq C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (5.11)
$\beta:=\left(\frac{2}{q_{1}}-\frac{1}{2}\right)\frac{\alpha}{\alpha+\delta},\quad\delta:=\frac{2}{q_{1}}-\frac{2}{q_{2}}.$
@@
    There exists a constant $C$ such that each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{2}(\xi_{n}))|\leq n$ and
@@
    Let $\phi\in V^{\prime}$ a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{2}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{Q}_{\Lambda_{2}(\xi_{n})}\langle\phi,u\rangle\right|\ \leq\ Cn^{-(2/q-1/2)}.$ (5.13)
@@ 6 Elliptic PDEs with affine inputs
Let $0<p\leq 2$. Let Assumption III hold. Let the assumptions of Lemma 6.1 hold for the spaces $W^{1}=V$ and $W^{r}$ with some $0<q_{1}\leq q_{r}<\infty$. For $\xi>1$, let $G(\xi)$ be the set defined by (2.7) for $\sigma_{1;{\boldsymbol{s}}}:=\beta_{1;{\boldsymbol{s}}}$ and $\sigma_{2;{\boldsymbol{s}}}:=\beta_{r;{\boldsymbol{s}}}$ as in (6.1). Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $\dim({\mathcal{V}}(G(\xi_{n}))\leq n$ and
@@
Let $1\leq p\leq\infty$. Let Assumption III hold. Let the assumptions of Lemma 6.1 hold for the spaces $W^{1}=V$ and $W^{r}$ with some $0<q_{1}\leq q_{r}<\infty$ with $q_{1}<2$. For $\xi>1$, let $G(\xi)$ be the set defined by in (3.5) for $\sigma_{1;{\boldsymbol{s}}}:=\beta_{1;{\boldsymbol{s}}}$ and $\sigma_{2;{\boldsymbol{s}}}:=\beta_{r;{\boldsymbol{s}}}$ as in (6.1). Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $\dim({\mathcal{V}}(G(\xi_{n}))\leq n$ and
@@
Let $1\leq p\leq\infty$. Let Assumption III hold. Let the assumptions of Lemma 6.1 hold for the spaces $W^{1}=V$ and $W^{r}$ with some $0<q_{1}\leq q_{r}<\infty$ with $q_{1}<2$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$. For $\xi>1$, let $G(\xi)$ be the set defined by (3.5) for $\sigma_{1;{\boldsymbol{s}}}:=\beta_{1;{\boldsymbol{s}}}$ and $\sigma_{2;{\boldsymbol{s}}}:=\beta_{r;{\boldsymbol{s}}}$ as in (6.1). Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that for the operator ${\mathcal{I}}_{G(\xi_{n})}:{\mathcal{L}}_{2}(W^{r})\to{\mathcal{V}}(G(\xi_{n})$, we have that $\dim{\mathcal{V}}(G(\xi_{n}))\leq n$ and
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
    $\left\|\int_{{\mathbb{I}}^{\infty}}u({\boldsymbol{y}})\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}u\right\|_{V}\leq C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (6.6)
@@
    $\left|\int_{{\mathbb{I}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}\langle\phi,u\rangle\right|\leq C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (6.7)
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
@@
    Let $\phi\in V^{\prime}$ be a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{I}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}(\xi_{n})}\langle\phi,u\rangle\right|\ \leq\ Cn^{-(2/q-1/2)}.$ (6.9)

### v14

@@ 5 Elliptic PDEs with lognormal inputs
where $S_{\Lambda}u:=\sum_{{\boldsymbol{s}}\in\Lambda}u_{\boldsymbol{s}}$ for $\Lambda\subset{\mathbb{F}}$, $k_{n}:=\lfloor\log_{2}\xi_{n}\rfloor$ if $\alpha\leq 1/q_{2}$, and $k_{n}:=\lfloor(\alpha q_{1})^{-1}\log_{2}\xi_{n}\rfloor$ if $\alpha>1/q_{2}$, and for $k\in{\mathbb{N}}_{0}$ and $\xi>1$,
@@
where $k_{\xi}:=\lfloor\log_{2}\xi\rfloor$ if $\alpha\leq 1/q_{2}$, and $k_{\xi}:=\lfloor(\alpha q_{1})^{-1}\log_{2}\xi\rfloor$ if $\alpha>1/q_{2}$, and $\Lambda_{k}(\xi)$ is defined as in (5.3), $\xi>1$. Moreover, $\Lambda_{k}(\xi)$ are downward closed sets, and consequently, the sequence $\big\{\Lambda_{k}(\xi)\big\}_{k=0}^{k_{\xi}}$ is nested in the inverse order, i.e., $\Lambda_{k^{\prime}}(\xi)\subset\Lambda_{k}(\xi)$ if $k^{\prime}>k$, and $\Lambda_{0}(\xi)$ is the largest and $\Lambda_{k_{\xi}}(\xi)$ is the smallest. Hence, the index set $G(\xi)$ can be constructed as in Algorithm 1.
@@ 28    $k\leftarrow k+1$
Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that for the interpolation operator ${\mathcal{I}}_{G(\xi_{n})}:{\mathcal{L}}_{2}^{\mathcal{E}}(W)\to{\mathcal{V}}(G(\xi_{n}))$, we have that $\dim{\mathcal{V}}(G(\xi_{n}))\leq n$ and
@@
The rate $\alpha$ corresponds to the spatial approximation of a single function in $W$ as given by (1.5). The rate $\beta$ is given by (3.21).
@@
Let $0<p\leq 2$. Under the hypothesis of Lemma 5.1 for the spaces $W^{1}=V$ with some $0<q_{1}=q<2$. For $\xi>1$, let $\Lambda(\xi)$ be the set defined by (3.42) for $\sigma_{\boldsymbol{s}}=\sigma_{1;{\boldsymbol{s}}}$ as in (5.1). Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda(\xi_{n}))|\leq n$ and
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
@@
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,{\rm d}\gamma({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}\langle\phi,v\rangle\right|\leq C\|\phi\|_{V^{\prime}}\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (5.7)
@@
$\beta:=\left(\frac{2}{q_{1}}-\frac{1}{2}\right)\frac{\alpha}{\alpha+\delta},\ \ \ \delta:=\frac{2}{q_{1}}-\frac{2}{q_{2}}.$
@@
where $k_{n}:=\lfloor\log_{2}\xi_{n}\rfloor$ if $\alpha\leq 1/q_{2}-1/2$, and $k_{n}:=\lfloor\vartheta\tau^{-1}\log_{2}\xi_{n}\rfloor$ if $\alpha>1/q_{2}-1/2$, and for $k\in{\mathbb{N}}_{0}$ and $\xi>1$,
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
@@
    Let $\phi\in V^{\prime}$ a bounded linear functional on $V$. There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}(\xi_{n})}\langle\phi,u\rangle\right|\ \leq\ C\|\phi\|_{V^{\prime}}n^{-(2/q-1/2)}.$ (5.9)
@@
Let Assumption I hold. Let the assumptions of Lemma 5.1 hold for the spaces $W^{1}=V$ and $W^{2}=W$ with some $0<q_{1}\leq q_{2}<\infty$ with $q_{1}<4$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$, and such that $Y_{m}$ is symmetric for every $m\in{\mathbb{N}}_{0}$. For $\xi>1$, let $G_{2}(\xi)$ be the set defined as in (LABEL:G_2(xi)) for $\sigma_{r;{\boldsymbol{s}}}$ as in (5.1), $r=1,2$. Then we have the following.
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G(\xi_{n}))\leq n$ and
@@
    Let $\phi\in V^{\prime}$ be a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G(\xi_{n}))\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{{\mathcal{Q}}}_{G_{2}(\xi_{n})}\langle\phi,v\rangle\right|\leq C\|\phi\|_{V^{\prime}}\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (5.11)
The rate $\alpha$ corresponds to the spatial approximation of a single function in $W$ as given by (1.5). The rate $\beta$ are given by
$\beta:=\left(\frac{2}{q_{1}}-\frac{1}{2}\right)\frac{\alpha}{\alpha+\delta},\ \ \ \delta:=\frac{2}{q_{1}}-\frac{2}{q_{2}}.$
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{2}(\xi_{n}))|\leq n$ and
@@
    Let $\phi\in V^{\prime}$ a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{2}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{Q}_{\Lambda_{2}(\xi_{n})}\langle\phi,u\rangle\right|\ \leq\ C\|\phi\|_{V^{\prime}}n^{-(2/q-1/2)}.$ (5.13)
@@ 6 Elliptic PDEs with affine inputs
Let $0<p\leq 2$. Let Assumption III hold. Let the assumptions of Lemma 6.1 hold for the spaces $W^{1}=V$ and $W^{r}$ with some $0<q_{1}\leq q_{r}<\infty$. For $\xi>1$, let $G(\xi)$ be the set defined by (2.7) for $\sigma_{1;{\boldsymbol{s}}}:=\beta_{1;{\boldsymbol{s}}}$ and $\sigma_{2;{\boldsymbol{s}}}:=\beta_{r;{\boldsymbol{s}}}$ as in (6.1). Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim({\mathcal{V}}(G(\xi_{n}))\leq n$ and
@@
Let $1\leq p\leq\infty$. Let Assumption III hold. Let the assumptions of Lemma 6.1 hold for the spaces $W^{1}=V$ and $W^{r}$ with some $0<q_{1}\leq q_{r}<\infty$ with $q_{1}<2$. For $\xi>1$, let $G(\xi)$ be the set defined by in (3.5) for $\sigma_{1;{\boldsymbol{s}}}:=\beta_{1;{\boldsymbol{s}}}$ and $\sigma_{2;{\boldsymbol{s}}}:=\beta_{r;{\boldsymbol{s}}}$ as in (6.1). Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim({\mathcal{V}}(G(\xi_{n}))\leq n$ and
@@
Let $1\leq p\leq\infty$. Let Assumption III hold. Let the assumptions of Lemma 6.1 hold for the spaces $W^{1}=V$ and $W^{r}$ with some $0<q_{1}\leq q_{r}<\infty$ with $q_{1}<2$. Assume that $(Y_{m})_{m\in{\mathbb{N}}_{0}}$ is a sequence satisfying the condition (3.15) for some positive numbers $\tau$ and $C$. For $\xi>1$, let $G(\xi)$ be the set defined by (3.5) for $\sigma_{1;{\boldsymbol{s}}}:=\beta_{1;{\boldsymbol{s}}}$ and $\sigma_{2;{\boldsymbol{s}}}:=\beta_{r;{\boldsymbol{s}}}$ as in (6.1). Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that for the operator ${\mathcal{I}}_{G(\xi_{n})}:{\mathcal{L}}_{2}(W^{r})\to{\mathcal{V}}(G(\xi_{n})$, we have that $\dim{\mathcal{V}}(G(\xi_{n}))\leq n$ and
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
    $\left\|\int_{{\mathbb{I}}^{\infty}}u({\boldsymbol{y}})\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}u\right\|_{V}\leq C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (6.6)
@@
    Let $\phi\in V^{\prime}$ be a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
    $\left|\int_{{\mathbb{I}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}\langle\phi,u\rangle\right|\leq C\|\phi\|_{V^{\prime}}\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (6.7)
@@
    There exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
@@
    Let $\phi\in V^{\prime}$ be a bounded linear functional on $V$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{I}}^{\infty}}\langle\phi,u({\boldsymbol{y}})\rangle\,\,{\rm d}\nu_{a,b}({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}(\xi_{n})}\langle\phi,u\rangle\right|\ \leq\ C\|\phi\|_{V^{\prime}}n^{-(2/q-1/2)}.$ (6.9)